# LoTLIP: Improving Language-Image Pre-training for Long Text Understanding

### v4

@@ Abstract
Understanding long text is of great demands in practice but beyond the reach of most language-image pre-training (LIP) models. In this work, we empirically confirm that the key reason causing such an issue is that the training images are usually paired with short captions, leaving certain tokens easily overshadowed by salient tokens. Towards this problem, our initial attempt is to relabel the data with long captions, however, directly learning with which may lead to performance degradation in understanding short text (e.g., in the image classification task). Then, after incorporating corner tokens to aggregate diverse textual information, we manage to help the model catch up to its original level of short text understanding yet greatly enhance its capability of long text understanding. We further look into whether the model can continuously benefit from longer captions and notice a clear trade-off between the performance and the efficiency. Finally, we validate the effectiveness of our approach using a self-constructed large-scale dataset, which consists of $100M$ long caption oriented text-image pairs. It is noteworthy that, on the task of long-text image retrieval, we beat the competitor using long captions with 11.1% improvement (i.e., from 72.62% to 83.72%). We will release the code, the model, and the new dataset to facilitate the reproducibility and further research. The project page is available at here.
@@ 1 Introduction
Understanding long texts is a vital skill in natural language processing (NLP), extensively utilized across various domains, including text generation [22, 30]. Building on this foundation, in the multi-modality field, some text-to-image works (e.g., DALLE-3 [1] and Pixel-art [4, 5]) employ pre-trained captioners (e.g., LLaVA [20]) to generate more accurate and detailed captions—in other words, long captions—for images. These long captions significantly reduce the generative model’s tendency to hallucinate, thereby enhancing the quality of text-image alignment. However, in the realm of language-image pre-training, scant research has been conducted on modeling long texts in a way that effectively aligns with image representations. Despite the lack of exploration, this is a critical problem with practical demands because multi-modality models are increasingly common in applications that require nuanced understanding of textual descriptions corresponding to images.
@@
To address these challenges, we have undertaken an extensive project to re-caption 100 million data with long captions, aiming to enrich the training environment for our models. This initiative allows us to explore the effects of increased text length in image-text pre-trained models and its impacts on model performance. Based on these experiments, we empirically confirm that the key reason causing such an issue is that the training images are usually paired with short captions, leaving certain tokens easily overshadowed by salient tokens. However, directly learning with long captions may improve the long-text understanding of image-text pre-trained models, but lead to performance degradation in understanding short texts (e.g., in the image classification task) as shown in Fig. 2. After integrating corner tokens to aggregate diverse textual information, we successfully enable the model to regain its original proficiency in understanding short texts while significantly improving its ability to comprehend long texts. We also explore whether the model can continue to benefit from longer captions and observe a clear trade-off between performance and efficiency. It is noteworthy that, on the task of long-text image retrieval, we beat Long-CLIP [35], a competitor using long captions for fine-tuning, with 11.1% improvement (i.e., from 72.62% to 83.72%).
@@ 2.1 Language-Image Pre-training
Recently, using language-image pre-trained models to do zero-shot prediction has attracted a lot of attention. CLIP [23] and ALIGN [14] demonstrate contrastive pre-trained models can learn rich visual-language correspondence knowledge from large-scale image-text pairs on the Internet and achieve good performance on zero-shot predictions, including image-text retrieval [18] and classification [9]. Following their success, various studies [31, 17, 16] have been devoted to improving image-text alignment. Among them, FILIP [31] focuses on fine-grained expressiveness between text tokens and image patches by modifying the training loss. While LiT [33] finds that apply contrastive-tuning to the pre-trained models with locked image encoder and unlocked text encoder can further improve the alignment. It is recognized that larger batch size brings better performance. For this reason, SigLIP [34] proposes to replace the softmax normalization among the standard contrastive loss with the sigmoid loss to scale up training batch size. In addition, LaCLIP [11], RLEG [36] and some other works [28, 13, 21] utilize multi-modality generative models to improve data quality for enhancing pre-training.
@@ 2.2 Long-text Understanding
Detailed long texts are necessary for many artificial intelligence tasks (e.g., human-computer interaction). Therefore, modeling and parsing long texts has become one of the most important research directions in natural language processing. Many studies in fields such as text generation [22, 30] have shown that advanced transformer structures have the ability to interpret long texts. However, in the field of language-image pre-training, research on using long-text descriptions of images to enhance multimodal representations is still very scarce. DreamLIP [37] utilizes multi-modality large language model to re-caption image data with detailed descriptions and then use them in language-image pre-training. In fact, during training, DreamLIP randomly extracts sub-captions from the detailed description for training and does not completely use all the information of the long text. One of the reasons why previous language-image pre-training rarely directly applied long texts for training is the text encoder of traditional CLIP [23] is restricted by the token number limit ($\leq 77$). Then, Long-CLIP [35] firstly introduces long captions into CLIP model to finetune, where the model is pre-trained on short-text-image datasets. The fine-tuning process equips the model with the ability to comprehend long texts, albeit at the expense of its proficiency in understanding shorter texts. In contrast, we incorporate long captions during the pre-training stage, which can not only enable the model to regain its original competency in interpreting short texts but also significantly improves its understanding of long texts. We further explored the trade-off between the benefit of long texts to the model and the training efficiency.
@@ 5.2 Ablation Studies
Exploring the Influence of Token Number Limitation
@@
Compare LoTLIP with other methods in pre-training with long texts
@@
Implementation of Corner Tokens
@@ 5.3 Main Results
We compare LoTLIP with the state-of-the-art methods on downstream tasks involving long texts and short texts. The experimental results are shown in Table 4 and Table 5, respectively. On 3M data scale, LoTLIP significantly improves the state-of-the-art methods on all tasks. Specifically, LoTLIP improves the second best method LiT by 29.99% on average over four long-text-image retrieval tasks. Moreover, LoTLIP improves the second competitor LiT by 6.58% and 10.98% on image classification task and short-text-image retrieval tasks, respectively. It proves that LoTLIP significantly enhances language-image model for understanding short and long captions by involving long captions in pre-training and incorporating corner tokens in text inputs. It is worth noting that LoTLIP trained with 100M data exceeds all state-of-the-art methods on long-text-image retrieval tasks, even though these methods are pre-trained on a larger scale of data. Concretely, LoTLIP (trained on 100M data) exceeds SigLIP (trained with 12B) by an average of 7.13% over four long-text-image retrieval tasks. Moreover, compared to Long-CLIP, which first uses long texts in CLIP for fine-tuning, LoTLIP improves averaged performance by 11.1% on long-text-image retrieval tasks.
@@ 6 Conclusion
In this work, we empirically confirm that a key issue arises because training images are typically paired with short captions, which can cause certain tokens to be overshadowed by more salient ones. To address this issue, our initial strategy involved relabeling the data with long captions. However, directly learning from these long captions might lead to degraded performance in tasks requiring an understanding of short text, such as image classification. Subsequently, by incorporating corner tokens to aggregate diverse textual information, we are able to help the model regain its original proficiency in understanding short texts while significantly enhancing its capability to comprehend long texts. We also investigated whether the model could continue to benefit from longer captions and observed a clear trade-off between performance and efficiency. Finally, we validated the effectiveness of our approach using a self-constructed large-scale dataset, which consists of 100 million long-caption-oriented text-image pairs. It is noteworthy that in the task of long-text image retrieval, our method outperforms the long caption-related competitor, Long-CLIP, with a significant improvement of 11.1%. We will release the code, the model, and the new dataset to facilitate the reproducibility and further research.

### v5

@@ Abstract
Understanding long text is of great demands in practice but beyond the reach of most language-image pre-training (LIP) models. In this work, we empirically confirm that the key reason causing such an issue is that the training images are usually paired with short captions, leaving certain tokens easily overshadowed by salient tokens. Towards this problem, our initial attempt is to relabel the data with long captions, however, directly learning with which may lead to performance degradation in understanding short text (e.g., in the image classification task). Then, after incorporating corner tokens to aggregate diverse textual information, we manage to help the model catch up to its original level of short text understanding yet greatly enhance its capability of long text understanding. We further look into whether the model can continuously benefit from longer captions and notice a clear trade-off between the performance and the efficiency. Finally, we validate the effectiveness of our approach using a self-constructed large-scale dataset, which consists of $100M$ long caption oriented text-image pairs. Our method demonstrates superior performance in long-text-image retrieval tasks. The project page is available here.
Corresponding authors.
@@ 1 Introduction
Understanding long texts plays a key role in natural language processing (NLP) field, e.g., text generation [23, 28, 33]. Inspired by these works, in the multi-modality field, some text-to-image generation works (e.g., DALLE-3 [1] and Pixel-art [4, 5]) employ pre-trained captioners (e.g., LLaVA [21]) to generate more accurate and detailed captions-in other words, long captions—for images. These long captions significantly reduce the generative model’s tendency to hallucinate, thereby enhancing the quality of text-image alignment. However, in the realm of language-image pre-training, scant research has been conducted on modeling long texts in a way that effectively aligns with image representations. Despite the lack of exploration, this is a critical problem with practical demands because multi-modality models are increasingly common in applications that require nuanced understanding of textual descriptions corresponding to images.
@@
To address these challenges, we have undertaken an extensive project to re-caption 100 million data with long captions, aiming to enrich the training environment for our models. This initiative allows us to explore the effects of increased text length in image-text pre-trained models and its impacts on model performance. Based on these experiments, we empirically confirm that the key reason causing such an issue is that the training images are usually paired with short captions, leaving certain tokens easily overshadowed by salient tokens. However, directly learning with long captions may improve the long-text understanding of image-text pre-trained models, but lead to performance degradation in understanding short texts (e.g., in the image classification task) as shown in Fig. 2. After integrating corner tokens to aggregate diverse textual information, we successfully enable the model to regain its original proficiency in understanding short texts while significantly improving its ability to comprehend long texts. We also explore whether the model can continue to benefit from longer captions and observe a clear trade-off between performance and efficiency. It is noteworthy that, on the task of long-text image retrieval, we beat Long-CLIP [39], a competitor using long captions for fine-tuning, with 4.2% improvement (i.e., from 79.52% to 83.72%).
@@ 2.1 Language-Image Pre-training
Recently, using language-image pre-trained models to do zero-shot prediction has attracted a lot of attention. CLIP [24] and ALIGN [14] demonstrate contrastive pre-trained models can learn rich visual-language correspondence knowledge from large-scale image-text pairs on the Internet and achieve good performance on zero-shot predictions, including image-text retrieval [18] and classification [9]. Following their success, various studies [16, 34, 36, 17] have been devoted to improving image-text alignment. Among them, FILIP [34] focuses on fine-grained expressiveness between text tokens and image patches by modifying the training loss. While LiT [37] finds that apply contrastive-tuning to the pre-trained models with locked image encoder and unlocked text encoder can further improve the alignment. It is recognized that larger batch size brings better performance. For this reason, SigLIP [38] proposes to replace the softmax normalization among the standard contrastive loss with the sigmoid loss to scale up training batch size. In addition, LaCLIP [11], RLEG [40] and some other works [19, 13, 22, 30] utilize multi-modality generative models to improve data quality for enhancing pre-training.
@@ 2.2 Long-text Understanding
Detailed long texts are necessary for many artificial intelligence tasks (e.g., human-computer interaction). Therefore, modeling and parsing long texts has become one of the most important research directions in natural language processing. Many studies in fields such as text generation [23, 32, 33] have shown that advanced transformer structures have the ability to interpret long texts. However, in the field of language-image pre-training, research on using long-text descriptions of images to enhance multimodal representations is still very scarce. DreamLIP [41] utilizes multi-modality large language model to re-caption image data with detailed descriptions and then use them in language-image pre-training. In fact, during training, DreamLIP randomly extracts sub-captions from the detailed description for training and does not completely use all the information of the long text. One of the reasons why previous language-image pre-training rarely directly applied long texts for training is the text encoder of traditional CLIP [24] is restricted by the token number limit ($\leq 77$). Then, Long-CLIP [39] firstly introduces long captions into CLIP model to finetune, where the model is pre-trained on short-text-image datasets. The fine-tuning process equips the model with the ability to comprehend long texts, albeit at the expense of its proficiency in understanding shorter texts. In contrast, we incorporate long captions during the pre-training stage, which can not only enable the model to regain its original competency in interpreting short texts but also significantly improves its understanding of long texts. We further explored the trade-off between the benefit of long texts to the model and the training efficiency.
@@ 5.2 Ablation Studies
Exploring the Influence of Token Number Limitation.
@@
Compare LoTLIP with Other Methods in Pre-training with Long Texts.
@@
Implementation of Corner Tokens.
@@ 5.3 Main Results
We compare LoTLIP with the state-of-the-art methods on downstream tasks involving long texts and short texts. The experimental results are shown in Table 4 and Table 5, respectively. On 3M data scale, LoTLIP significantly improves the state-of-the-art methods on all tasks. Specifically, LoTLIP improves the second best method LiT by 29.99% on average over four long-text-image retrieval tasks. Moreover, LoTLIP improves the second competitor LiT by 6.58% and 10.98% on image classification task and short-text-image retrieval tasks, respectively. It proves that LoTLIP significantly enhances language-image model for understanding short and long captions by involving long captions in pre-training and incorporating corner tokens in text inputs. It is worth noting that LoTLIP trained with 100M data exceeds all state-of-the-art methods on long-text-image retrieval tasks, even though these methods are pre-trained on a larger scale of data. Concretely, LoTLIP (trained on 100M data) exceeds SigLIP (trained with 12B) by an average of 7.13% over four long-text-image retrieval tasks. Moreover, compared to Long-CLIP, which first uses long texts in CLIP for fine-tuning, LoTLIP improves averaged performance by 4.2% on long-text-image retrieval tasks.
@@ 6 Conclusion
In this work, we empirically confirm that a key issue arises because training images are typically paired with short captions, which can cause certain tokens to be overshadowed by more salient ones. To address this issue, our initial strategy involved relabeling the data with long captions. However, directly learning from these long captions might lead to degraded performance in tasks requiring an understanding of short text, such as image classification. Subsequently, by incorporating corner tokens to aggregate diverse textual information, we are able to help the model regain its original proficiency in understanding short texts while significantly enhancing its capability to comprehend long texts. We also investigated whether the model could continue to benefit from longer captions and observed a clear trade-off between performance and efficiency. Finally, we validated the effectiveness of our approach using a self-constructed large-scale dataset, which consists of 100 million long-caption-oriented text-image pairs. It is noteworthy that in the task of long-text image retrieval, our method outperforms the long caption-related competitor, Long-CLIP, with a significant improvement of 4.2%.
7 Acknowledgments
This work was supported by National Natural Science Foundation of China (NSFC) under Grants 62225207, and Zhejiang Provincial Natural Science Foundation of China under Grants LD24F020011.